A 40 ohm resistor and a 60 ohm resistor are connected in series to a 12 volt battery.


A. What is the voltage difference across each resistor?
Answer #1

Current = 12/(40+60) = 12/100 = 0.12 A

A)

Voltage accross 40 ohm = I*R = 0.12*40 = 4.8 V

Voltage accross 60 ohm = I*R = 0.12*60 = 7.2 V
Answer #2

Req = 40 + 60 = 100 Ohm


i = V / Req = 12 / 100 = 0.12 A


A)

Voltage across 40 => 0.12*40 = 4.8 V

Voltage across 60 => 0.12*60 = 7.2 V
Answer #3

If resistors are connected in series them same current flows

Given

R1=40 ohms R2=60 ohms V=12 volts

Now total voltage=V=v1+v2

Total resistance=R=r1+r2=40+60=100 (since they are in series)

Now apply ohms law

V=IR substituting values we get

12=I*100

I=12/100 amperes(total current)

Now v1= voltage across 40 ohms resistor =IR1=(12/100)*40=4.8 volts

V2=voltage across 60 ohms resistor=IR2=(12/100)*60=7.2 volts